# IRIS: An Immersive Robot Interaction System

### v2

@@ III-A1 Node Communication Protocol
The IRIS system operates across simulation and/or sensor-processing computers, multiple XR headsets, and other monitoring and control programs, requiring a robust and reliable network connection between them. All devices are part of the same subnet, with all the devices connected via Wi-Fi or cable. Inspired by the Robot Operating System (ROS [50]), However, instead of using a single host IRIS leverages a local network across multiple hosts, while using both Request-Response and Publish-Subscribe patterns for data transmission. This protocol follows a master-node architecture with multiple nodes, where the simulation PC serves as the master node, and the XR headsets and other devices act as XR nodes. Communication is achieved through a combination of UDP sockets and ZeroMQ (ZMQ).
@@
III-A2 Unified Scene Specification
@@
III-A3 Extendable and Flexible Framework Support
@@
IRIS provides a user-friendly API. For each environment or framework, a single line of code suffices to visualize and update the simulation in the XR headset. Here is a short example of how to use it in the MuJoCo Simulation, where the important line is marked in bold:
1# import scenepub
2from scenepub.sim.mj_publisher import MujocoPublisher
3# define the mujoco environment
4model = mujoco.MjModel.from_xml_path(xml_path)
5data = mujoco.MjData(model)
6# define the ScenePublisher for mujoco
7publisher = MujocoPublisher(model, data, host)
8# run simulation
9while True:
10 # run simulation step logic
11 pass
@@
III-A4 Real Scene Loading
@@
III-A5 Interaction Data Collection
IRIS provides user-friendly access to interaction data through the ScenePublisher, which supports both Meta Quest 3 and HoloLens 2. Users only need to create a new device instance and assign a name to the connection. The ScenePublisher then waits automatically for the device to launch. Once the device is launched, the instance receives messages from the XR headset, enabling users to access various types of interaction data through different methods. This process makes it straightforward to retrieve data such as hand tracking, motion controller inputs, and other relevant interaction information. The interaction data accessible through the ScenePublisher is listed in Tab. II, and the example code can be found in the Appendix. B-G
Feature HoloLens 2 Meta Quest 3
Hand Tracking
Gaze Tracking
Motion Controller
Head Tracking
TABLE II: Feature comparison between HoloLens 2 and Meta Quest 3. IRIS supports the transmission of interaction data from headsets to the simulation environment, with easy access to this data through ScenePublisher.
III-A6 Spatial Anchor
A spatial anchor serves as a reference point in a 3D environment to accurately position virtual objects within physical space. It enables virtual objects to maintain their position, orientation, and alignment, even as users move around or leave and return to the environment. IRIS utilizes either QR codes or motion controllers (currently only implemented for the Meta Quest 3) as spatial anchors. This approach allows for the alignment of augmented scenes across multiple headsets in the real world, making it appear as though all users are sharing the same scene. Additionally, this method can synchronize a virtual robot with its real-world counterpart in the Kinesthetic Teaching (III-B1) interface and facilitate multiple-collector view alignment. Fig. 5 shows the alignment between real robot and virtual robot.
Figure 5: This graph shows the alignment between real robot and virtual robot by a spatial anchor, which is used for controlling virtual robot by real world Kinesthetic Teaching III-B1.
III-A7 Headset Compatibility
IRIS implements an XR application using Unity, featuring essential capabilities such as node communication protocols, scene construction based on scene specifications, and synchronization with remote simulations. The application can be directly deployed to other headset platforms using the Unity deployment pipeline, showcasing IRIS’s Cross-Platform capability. For input data handling (reading and sending), the application requires the use of platform-specific APIs, making it impractical to rely on a generalized framework. This approach separates the visualization and interaction components, minimizing the effort needed to transfer IRIS codebase to new platforms.
III-B Intuitive Robot Control Interface
III-B1 Kinesthetic Teaching
Kinesthetic teaching is an intuitive interface that allows users to control robots by physically moving the real robot [62, 54, 26]. The real robot transmits joint positions and velocities in real time to the controlled robot, which can be either a virtual or another physical robot. In both real-world and simulation data collection, the key aspect of kinesthetic teaching is ensuring alignment between the real robot and its virtual counterpart in the XR headsets. By utilizing Spatial Anchors (III-A6), the virtual robot in the XR headsets can be perfectly aligned with the real robot, which provide users with an intuitive and immersive experience.
III-B2 Motion Controller
III-C Affiliated Monitor Tools
The extensibility of IRIS opens up numerous possibilities for creating new applications. IRIS includes a web-based monitoring tool for managing all XR headsets. This tool allows users to easily start and stop alignment processes, as well as rename devices. Additionally, the tool supports real-time scene visualization using three.js [3], using the unified scene specification. An example screenshot is shown in Fig. 7, with further technical details provided in the appendix.
IV System Application
IV-A Robot Data Collection for Benchmarks
Through its flexible framework design, IRIS supports four simulators and various robot manipulation benchmarks. Based on the interfaces introduced in Sec. III-A5, IRIS already provides example controllers for various benchmarks and frameworks, such as robosuite, LIBERO, RoboCasa, mink, Metaworld, Fancy Gym, and so on. Robots from all the frameworks can be controlled by Motion Controller or Kinesthetic Teaching. Fig. 1 shows the robots controlled by IRIS including Franka Panda, Aloha 2, Barrett Wam Arm, UR5e, iiwa14, Boston Dynamics Spot, Unitree H1, and more.
IV-B Collaborative Manipulation
Figure 8: Collaborative manipulation in the simulation via XR. The left image shows the collaborative manipulation for hand over a hammer between two Franka Panda robots by Kinesthetic Teaching, and the right image shows that collaborative manipulation for hand over a red board between two Aloha 2 Arms by Motion Controller.
Collaborative manipulation, where multiple users provide demonstrations simultaneously and/or interactively, plays a crucial role in the advancement of human-robot systems [58]. Collecting demonstration data for such tasks has been a significant challenge as it requires smooth communication and synchronization between human participants, robots, and the virtual environment. Previous work on manipulation involving multiple humans often facilitates collaboration through multiple screens [49], which lacks the immersive experience provided by XR. Moreover, typically only one person is in control of the demonstrations while others are limited to observing or monitoring [55]. By leveraging the node communication protocol, IRIS allows for easy integration of new devices and interfaces for controlling multiple robots in a shared scene. Specifically, IRIS supports the addition of new XR headsets with minimal effort, allowing more users to dynamically join the collaborative environment. Fig. 8 shows collaborative manipulation for handover task.
IV-C Interaction with Robot Policies in Simulation
Simulation has long been an essential tool for training robot agents, especially for reinforcement learning (RL) policies. Its advantages – such as safety, scalability, and cost-efficiency – make it widely used in RL training. Previous works on interactive learning have utilized interfaces such as keyboards [32], 3D mouse [31, 29] or smartphone [33]. However, these methods are insufficient for tasks that require multiple viewpoints or complex motions. IRIS delivers an immersive experience, providing the appearance of "stepping into" the simulation. Its interaction API (III-A5) allows users to effortlessly move and manipulate objects within the simulation, mirroring real-world interactions. This capability enables IRIS to serve as a powerful tool for highly interactive tasks, such as competitive sports, where real-time responsiveness and precise control are essential.
(a) Interact with RL Agent in the first-person view (left) and simulation (right)
(b) Interact with RL Agent in the third-person view
Figure 9: Playing table tennis with RL agent in Fancy Gym environment, the RL agent policy is trained with Deep Black-Box Reinforcement Learning (BBRL) [45]
To demonstrate the capabilities of IRIS, we designed an experiment in which a participant played table tennis against an episodic RL agent trained with Deep Black-Box Reinforcement Learning (BBRL) [45] exclusively in simulation. The environment was adapted from Fancy Gym [44]. Table tennis is a challenging task that requires dynamic perspective shifts and precise racket control. Using a motion controller and XR headset, the participant learned to successfully return the ball the RL agent served within a few minutes. This success highlights the potential of IRIS to facilitate high-quality interactive data collection, which can further enhance the training and performance of RL agents.
IV-D Real World Teleoperation and Data Collection
Real robot data plays a crucial role in conducting real-world experiments. To minimize physical obstruction from humans, tele-operation using XR (Extended Reality) is commonly employed for collecting such data. Some approaches, such as [23], utilize cameras to stream videos to XR headsets. However, this method lacks depth perception. Other works, like [14], employ movable platforms to track the movement of users’ heads, allowing for active scene observation by adjusting the viewing perspective. While effective, this approach requires significant effort and resources to install and maintain the necessary hardware.
To overcome these limitations, we integrate a point cloud-based XR tele-operation system into our tele-operation framework, ensuring both immersion and interactivity for data collection. An ORBBEC Femto Bolt depth camera captures real-time depth and RGB data, and a QR code on the robot’s end effector enables camera-to-robot base transformation, aligning point clouds with the robot’s frame. A GPU-accelerated C++ pipeline processes (see Appendix B-E2) point clouds at 30Hz (~2 million points per frame), applying voxel grid downsampling to reduce computational load and latency while maintaining real-time performance. Before downsampling, unnecessary regions are cropped, retaining only relevant portions of the workspace. The processed point clouds are transmitted to a main process, which forwards them to a Meta Quest 3 headset for real-time visualization and interaction.
Figure 10: Real-world teleoperation system integrating augmented reality for intuitive robot control. The setup includes two Franka Emika Panda robots: a Leader Robot (cyan) controlled by a Meta Quest 3 user and a Follower Robot (blue) mirroring its movements. An ORBBEC Femto Bolt depth camera (yellow) captures the scene for real-time point cloud visualization (green). The top-left and top-right images show the augmented view, reconstructing objects (e.g., red and yellow cups) in XR. The bottom image presents the physical setup, where the Leader Robot is directly controlled while the Follower Robot autonomously replicates actions, enabling precise, immersive remote operation.
Our teleoperation framework features two Franka Emika Panda [22] robots: a leader controlled by a human in zero-torque mode and a follower mirroring it’s movements, including gripper actions. This setup enables seamless remote manipulation while eliminating the need for human presence in the physical workspace. In future implementations, additional cameras can be incorporated, and point cloud fusion via Iterative Closest Point (ICP) will ensure precise alignment across multiple views. The system architecture, including point cloud processing, XR integration, and teleoperation, is illustrated in Figure 10. By integrating real-time point cloud visualization with XR-based teleoperation, our system enhances efficiency, accuracy, and scalability, enabling obstruction-free data collection for human-robot interaction studies.
V Experiment
This section evaluates IRIS’s capability to create demonstrations, focusing on the efficiency and intuitiveness of data collection. It is followed by a performance profile analysis.
V-A Data Collection Evaluation in Simulation
To assess the data collection efficiency of IRIS, a user study was conducted by using tasks from the LIBERO benchmark [29], which represents diverse types of movements. Four representative tasks (see Figure 11) were selected in the dimension of translation, rotation, and compound movement:
    Task 1: close the microwave
    Task 2: turn off the stove
    Task 3: pick up the book in the middle and place it on the cabinet shelf
    Task 4: turn on the stove and put the frying pan on it.
Refer to caption Refer to caption Refer to caption Refer to caption
(a1) (b1) (c1) (d1)
Refer to caption Refer to caption Refer to caption Refer to caption
(a2) (b2) (c2) (d2)
Interface Task 1 Task 2 Task 3 Task 4
Keyboard $0.90$ $0.725$ $0.750$ $0.500$
3D Mouse $1.00$ $0.950$ $0.375$ $0.900$
Kinesthetic Teaching (Ours) $1.00$ $1.00$ $0.975$ $0.950$
Motion Controller (Ours) $1.00$ $0.900$ $0.975$ $1.00$
TABLE III: Success rate of four different interfaces (percentage). The baselines are Keyboard and 3D Mouse from LIBERO. The results show that both Kinesthetic Teaching and the Motion Controller from IRIS consistently achieve a high success rate, nearing 100%, across all four task.
Figure 12: This graph shows the average task completion time (in seconds) for each interface across tasks. The KT and MC interfaces consistently perform more efficiently, while the Keyboard and 3D Mouse interfaces result in longer completion times, particularly on more complex tasks.
For the objective metrics, Table III presents the success rates for each interface across the tasks. A trial was considered unsuccessful if the participant failed to complete the task or exceeded the time limit. The time limits were set based on task difficulty: 20 seconds for Task 1 and Task 2, 30 seconds for Task 3, and 40 seconds for Task 4. The data shows a success rate of over $90\%$ across all four tasks when using the KT and MC interfaces from IRIS. In contrast, the Keyboard and 3D Mouse methods from LIBERO often resulted in failures. For example, the 3D Mouse interface achieved only a $37.5\%$ success rate on task 3. Figure 12 shows the average time consumed for each task across four interfaces. The KT and MC interfaces consistently demonstrate lower task completion times, indicating higher efficiency. In contrast, the Keyboard and 3D Mouse interfaces show significantly higher completion times, particularly for Task 3 and Task 4, where the 3D Mouse method approaches or exceeds the task’s time limit. These results align with the observed lower success rates for these interfaces, highlighting their inefficiency in time-critical tasks.
Figure 13: Subjective evaluation scores for usefulness, experience, intuitiveness, and efficiency across four interfaces. The KT and MC interfaces perform favorably in all categories, while the Keyboard and 3D Mouse interfaces receive lower ratings, particularly in intuitiveness and efficiency.
Figure 13 compares subjective scores for four interfaces based on usefulness, experience, intuitiveness, and efficiency. The KT and MC interfaces consistently receive high scores across all criteria, indicating positive user perception and ease of use. In contrast, the Keyboard and 3D Mouse interfaces receive significantly lower ratings, particularly in intuitiveness and efficiency, reflecting the participants’ difficulties in using these methods to control the robot.
The results of this user study show that IRIS received higher scores than the baseline interfaces across both objective and subjective metrics. This indicates that the system offers a more intuitive and efficient approach for data collection.
V-B Performance Analysis
The performance analysis focuses on two key aspects: network latency and headset FPS (frames per second).
Since IRIS uses asynchronous bidirectional data transfer, network latency is low, averaging around 20-30 ms. Transmission bandwidth depends on Wi-Fi capacity. Even for large scenes from RoboCasa, which contain over 200 MB of compressed assets, it takes no more than 5 seconds to transfer and generate a full scene with more than 300 objects. For real robot teleoperation, the system handles point cloud data efficiently, achieving a transmission speed of 10,000 points at 60 Hz—exceeding the camera’s frame rate.
VI Limitations
Based on our development and usage experience, IRIS has three main limitations. First, the visualization in XR headsets is limited to basic RGB textures for materials, making robots from different simulators appear visually identical. Second, IRIS currently only supports rigid objects, at least deformable object have not yet been fully explored. Finally, IRIS has only been tested on Meta Quest 3 and HoloLens 2, and further evaluation on additional devices would be beneficial.
VII Conclusion
In this work, we introduced IRIS (Immersive Robot Interaction System), an innovative framework that seamlessly integrates Extended Reality (XR) technologies with robotics data collection. IRIS addresses key challenges in reproducibility and reusability that are common in current XR-based systems. Its flexible, extendable design supports multiple simulators, benchmarks, real-world applications, and multi-user use cases. User studies demonstrate that IRIS outperforms previous data collection methods, positioning it as a promising solution for future data collection pipelines. As an open-source project, IRIS codebase promotes further research and adaptation across diverse use cases and hardware platforms.
Appendix A Appendix Title
Appendix B Technical Details
B-A Mujoco
@@
1def mj2unity_pos(pos: List[float]) -> List[float]:
2 return [-pos[1], pos[2], pos[0]]
3
4def mj2unity_quat(quat: List[float]) -> List[float]:
5 return [quat[2], -quat[3], -quat[1], quat[0]]
@@
1class MujocoPublisher(ScenePublisher):
2
3 def __init__(
4 self,
5 mj_model,
6 mj_data,
7 host: str = "127.0.0.1",
8 no_rendered_objects: Optional[List[str]] = None,
9 no_tracked_objects: Optional[List[str]] = None,
10 visible_geoms_groups: Optional[List[int]] = None,
11 ) -> None:
B-B IsaacSim
@@
The scenes in IsaacSim are organized in a Universal Scene Description (USD) format, and data can be accessed directly with the OpenUSD API [53]. The scene hierarchy consists of transformations (Xform), meshes, articulations (joints) among other elements. Figure 14 illustrates a sample scene hierarchy.
@@
1# compute local transforms of the current prim
2trans, rot, scale = self.compute_local_trans(root)
3
4# parse material and geometry
5mat_info = self.parse_prim_material(prim=root)
6self.parse_prim_geometries(
7 prim=root,
8 prim_path=prim_path,
9 sim_obj=sim_object,
10 mat_info=mat_info or inherited_material,
11)
12
13# parse children of the current prim
14for child in root.GetChildren():
15 if obj := self.parse_prim_tree(
16 root=child,
17 parent_path=prim_path,
18 inherited_material=mat_info or inherited_material,
19 ):
20 sim_object.children.append(obj)
@@
1from pxr import Usd
2from omni.isaac.core.prims import XFormPrim
3
4def compute_world_trans(self, prim: Usd.Prim):
5 prim = XFormPrim(str(prim.GetPath()))
6 assert prim.is_valid()
7
8 pos, quat = prim.get_world_pose()
9 scale = prim.get_world_scale()
10
11 return (
12 pos.cpu().numpy(),
13 quat_to_rot_matrix(quat.cpu().numpy()),
14 scale.cpu().numpy(),
15 )
B-C CoppeliaSim
CoppeliaSim, formerly known as V-REP, is a versatile and widely used robot simulation software that supports various physics engine backbones. Scenes in CoppeliaSim are constructed using a tree structure, which includes objects such as visual shapes, dynamic shapes, and joints, as illustrated in Figure 15.
@@ B-C CoppeliaSim
1from coppeliasim_zmqremoteapi_client import RemoteAPIClient
2client = RemoteAPIClient()
3sim = client.require(’sim’)
4
5list_vertices, list_indices, list_normals = [], [], []
6
7objects_id_list = sim.getObjectsInTree(sim.handle_scene,
8 sim.handle_all, 0)
9
10for idx in objects_id_list:
11 if sim.getObjectType(idx) == sim.sceneobject_shape:
12 # We assume all the shapes are primary shapes
13 vertices, indices, normals = sim.getShapeMesh(idx)
14
15 list_vertices.append(vertices)
16 list_indices.append(indices)
17 list_list_normals.append(normals)
@@
1from pyrep.backend.sim import simGetObjectsInTree
2from pyrep.backend.sim import simGetObjectType
3from pyrep.backend.sim import simGetShapeMesh
4from pyrep.backend.simConst import sim_handle_scene
5from pyrep.backend.simConst import sim_handle_all
6from pyrep.backend.simConst import sim_object_shape_type
7
8list_vertices, list_indices, list_normals = [], [], []
9
10objects_id_list = simGetObjectsInTree(sim_handle_scene,
11 sim_handle_all, 0)
12
13for idx in objects_id_list:
14 if simGetObjectType(idx) == sim_object_shape_type:
15 # We assume all the shapes are primary shapes
16 vertices, indices, normals = simGetShapeMesh(idx)
17
18 list_vertices.append(vertices)
19 list_indices.append(indices)
20 list_list_normals.append(normals)
@@ B-D Genesis
Each rigid entity (RigidEntity) from Genesis is processed to create a SimObject. The position and rotation of the entity are retrieved and transformed is similart to Mujoco (B-A) to ensure compatibility with Unity’s coordinate system. If the Genesis scene is already built, the function pulls actual position and rotation data; otherwise, it defaults to identity transforms.
@@ B-E Real World
B-E1 Camera-to-Robot Base Transformation

### v3

@@ A.1 Node Communication Protocol
The IRIS system operates across simulation and/or sensor-processing computers, multiple XR headsets, and other monitoring and control programs, requiring a robust and reliable network connection between them. All devices are part of the same subnet, with all the devices connected via Wi-Fi or cable. Different from Robot Operating System (ROS [40]), IRIS leverages a local network across multiple hosts instead of using a single host, while using both Request-Response and Publish-Subscribe patterns for data transmission. This protocol follows a master-node architecture, where the simulation PC serves as the master node, and the XR headsets and other devices act as XR nodes. Communication is achieved through a combination of UDP sockets and ZeroMQ (ZMQ).
@@
A.2 Unified Scene Specification
@@
A.3 Extendable and Flexible Framework Support
@@
IRIS provides a user-friendly API. For each environment or framework, a single line of code suffices to visualize and update the simulation in the XR headset. Here is a short example of how to use it in the MuJoCo Simulation:
# import scenepub
from scenepub.sim.mj_publisher import MujocoPublisher
# define the mujoco environment
model = mujoco.MjModel.from_xml_path(xml_path)
data = mujoco.MjData(model)
# define the ScenePublisher for mujoco; only this line needs to be added
publisher = MujocoPublisher(model, data, host)
# run simulation
while True:
# run simulation step logic
pass
@@
A.4 Real Scene Loading
@@
Appendix B Simulator Support and Real-world Data Collection
B.1 Mujoco
@@
def mj2unity_pos(pos: List[float]) -> List[float]:
return [-pos[1], pos[2], pos[0]]
def mj2unity_quat(quat: List[float]) -> List[float]:
return [quat[2], -quat[3], -quat[1], quat[0]]
@@
class MujocoPublisher(ScenePublisher):
def __init__(
self,
mj_model,
mj_data,
host: str = "127.0.0.1",
no_rendered_objects: Optional[List[str]] = None,
no_tracked_objects: Optional[List[str]] = None,
visible_geoms_groups: Optional[List[int]] = None,
) -> None:
B.2 IsaacSim
@@
The scenes in IsaacSim are organized in a Universal Scene Description (USD) format, and data can be accessed directly with the OpenUSD API [68]. The scene hierarchy consists of transformations (Xform), meshes, articulations (joints) among other elements. Fig. 12 illustrates a sample scene hierarchy.
@@
# compute local transforms of the current prim
trans, rot, scale = self.compute_local_trans(root)
# parse material and geometry
mat_info = self.parse_prim_material(prim=root)
self.parse_prim_geometries(
prim=root,
prim_path=prim_path,
sim_obj=sim_object,
mat_info=mat_info or inherited_material,
# parse children of the current prim
for child in root.GetChildren():
if obj := self.parse_prim_tree(
root=child,
parent_path=prim_path,
inherited_material=mat_info or inherited_material,
sim_object.children.append(obj)
@@
from pxr import Usd
from omni.isaac.core.prims import XFormPrim
def compute_world_trans(self, prim: Usd.Prim):
prim = XFormPrim(str(prim.GetPath()))
assert prim.is_valid()
pos, quat = prim.get_world_pose()
scale = prim.get_world_scale()
return (
pos.cpu().numpy(),
quat_to_rot_matrix(quat.cpu().numpy()),
scale.cpu().numpy(),
B.3 CoppeliaSim
CoppeliaSim, formerly known as V-REP, is a versatile and widely used robot simulation software that supports various physics engine backbones. Scenes in CoppeliaSim are constructed using a tree structure, which includes objects such as visual shapes, dynamic shapes, and joints, as illustrated in Fig. 13.
@@ B.3 CoppeliaSim
from coppeliasim_zmqremoteapi_client import RemoteAPIClient
client = RemoteAPIClient()
sim = client.require(’sim’)
list_vertices, list_indices, list_normals = [], [], []
objects_id_list = sim.getObjectsInTree(sim.handle_scene,
sim.handle_all, 0)
for idx in objects_id_list:
if sim.getObjectType(idx) == sim.sceneobject_shape:
# We assume all the shapes are primary shapes
vertices, indices, normals = sim.getShapeMesh(idx)
list_vertices.append(vertices)
list_indices.append(indices)
list_list_normals.append(normals)
@@
from pyrep.backend.sim import simGetObjectsInTree
from pyrep.backend.sim import simGetObjectType
from pyrep.backend.sim import simGetShapeMesh
from pyrep.backend.simConst import sim_handle_scene
from pyrep.backend.simConst import sim_handle_all
from pyrep.backend.simConst import sim_object_shape_type
list_vertices, list_indices, list_normals = [], [], []
objects_id_list = simGetObjectsInTree(sim_handle_scene,
sim_handle_all, 0)
for idx in objects_id_list:
if simGetObjectType(idx) == sim_object_shape_type:
# We assume all the shapes are primary shapes
vertices, indices, normals = simGetShapeMesh(idx)
list_vertices.append(vertices)
list_indices.append(indices)
list_list_normals.append(normals)
@@ B.4 Genesis
Each rigid entity (RigidEntity) from Genesis is processed to create a SimObject. The position and rotation of the entity are retrieved and transformed is similart to Mujoco (B.1) to ensure compatibility with Unity’s coordinate system. If the Genesis scene is already built, the function pulls actual position and rotation data; otherwise, it defaults to identity transforms.
@@ B.5 Real World
B.5.1 Camera-to-Robot Base Transformation